# JailbreakLens: Interpreting Jailbreak Mechanism in the Lens of Representation and Circuit

### v1

@@ 6 Discussion
Limitations. While this work offers valuable insights into the mechanisms behind jailbreak attacks on LLMs, there are several limitations worth noting. Firstly, due to computational resource constraints, the experiments are limited to models with a maximum size of 13B parameters. Larger models, such as those with 60B parameters or more, are not explored. Since larger models may exhibit different vulnerabilities or resistance to jailbreak attacks, future studies should extend this research to include larger-scale models to determine whether similar attack patterns and mitigation strategies apply.
@@
Secondly, we focus on established categorization of jailbreak strategies and select one or two representative methods from each category for our analysis. Some of the more recent jailbreak methods are not included in this study. However, our findings suggest that the underlying mechanisms driving the success of different jailbreak methods within the same category are fundamentally similar, implying that the insights gained from our research may be broadly applicable.
@@
Future direction. Future work will build on the insights gained from this research to design more robust and interpretable safety mechanisms for LLMs. Specifically, by leveraging the understanding of how jailbreak methods manipulate model components, such as key attention heads and MLP layers, we can develop more generalizable and resilient safeguards. These defenses could target the specific areas of vulnerability identified, such as the suppression of refusal signals and the amplification of affirmation signals.
7 Conclusion
@@
In this paper, we propose a dual-perspective framework, named JailbreakLens, which explains the jailbreak mechanism from both representations and circuit levels. Through an in-depth analysis of seven jailbreak methods on the Llama and Vicuna models, we find jailbreak prompts amplify components that reinforce affirmative responses while suppressing those that produce refusal. Although this manipulation shifts model representations toward safe clusters to deceive the LLM, leading it to provide detailed responses instead of refusals, it still produce abnormal activation which can be caught in the circuit analysis. Additionally, we found that increasing the model scale does not significantly enhance resistance to jailbreaks, and fine-tuning has a limited effect on the model’s inner safety circuit. These insights deepen our understanding of how jailbreaks succeed and provide valuable guidance for developing stronger defense mechanisms.
Appendix A Details of Jailbreak Attacks
@@
TABLE V: Attack success rate of harmful prompts and prompts with different jailbreak strategies.
The attack success rate of the jailbreak strategies considered in this paper on the Advbench dataset is summarized in Tab. V. Here are examples of each jailbreak method targeted Llama2-7b used for evaluation.
@@
Appendix B Supplementary Experimental Results
B.1 Supplementary Results of Circuit Analysis
The contributions to the refusal score of all attention heads in all models we analyzed are shown in Fig. 7. For the models For Vicuna1.5-7b, the location of affirmation and refusal signal attention heads is consistent with Llama2-7b, i.e., L26H4 and L21H14. For Llama2-13b and Vicuna1.5-13b, the location of affirmation and refusal signal attention heads is the same, i.e., L37H37 and L31H35. Moreover, the results of contributions to the refusal score of MLP layers are shown in Fig. 8.
The results of how signal heads and signal MLP layers are activated with jailbreak prompts are summarized in Tab. VI and Fig. 9, respectively.
(a) Harmful, Llama2-7b
(b) Safe, Llama2-7b
(c) Harmful, Vicuna1.5-7b
(d) Safe, Vicuna1.5-7b
(e) Harmful, Llama2-13b
(f) Safe, Llama2-13b
(g) Harmful, Vicuna1.5-13b
(h) Safe, Vicuna1.5-13b
Figure 7: Attention Heads attribution of refusal score in Llama2-7b, and Vicuna1.5-7b, and Llama2-13b, and Vicuna1.5-13b, respectively.
(a) Llama2-7b
(b) Llama2-13b
(c) Vicuna1.5-7b
(d) Vicuna1.5-13b
Figure 8: MLP layers attribution of refusal score on harmful and safe prompts in Llama2-7b, and Vicuna1.5-7b, and Llama2-13b, and Vicuna1.5-13b, respectively.
(a) Llama2-7b, jailbreaks
(b) Llama2-13b, jailbreaks
(c) Vicuna1.5-7b, jailbreaks
(d) Vicuna1.5-13b, jailbreaks
Figure 9: MLP layers attribution of refusal score on jailbreak prompts in Llama2-7b, and Vicuna1.5-7b, and Llama2-13b, and Vicuna1.5-13b, respectively.
Prompt Method Llama2-7b Llama2-13b Vicuna1.5-7b Vicuna1.5-13b
$S_{+}$ $S_{-}$ $S_{+}$ $S_{-}$ $S_{+}$ $S_{-}$ $S_{+}$ $S_{-}$
Baseline Safe 1.0 -0.0760 1.0 0.1873 1.0 0.4343 1.0 0.1097
Harmful -0.0241 1.0 -0.1457 1.0 0.1562 1.0 -0.0789 1.0
Jailbreak GCG 0.9819 1.0 1.0 0.8515 0.6021 1.0 1.0 0.7339
ReNellm 1.0 0.3816 1.0 0.0442 0.7485 1.0 1.0 0.2915
PAIR 0.9810 1.0 0.5479 1.0 0.5170 0.6897 0.3374 1.0
DeepInception 1.0 -0.0997 0.9956 0.6971 1.0 0.7607 0.8672 0.7089
Autodan 0.0202 0.7573 0.0281 0.6594 0.4675 0.6523 0.5013 0.6127
GPTFuzz 0.0702 0.2817 0.3298 0.1600 0.4643 1.0 0.4819 0.6724
CodeChameleon 1.0 0.4816 0.9116 0.2723 0.4983 1.0 0.5439 0.8726
TABLE VI: Refusal attribution of each attention head Activation of signal attention heads, i.e., $S_{+}$ denotes affirmation signal head and $S_{-}$ denotes refusal signal head, of Llama-2, and Vicuna-1.5 series’ models on different jailbreak methods.
B.2 Supplementary Results of Evolution Analysis
@@
Figure 10: Dynamics of key components activation and representation on different jailbreak prompts in Llama2-7b.

### v2

@@ 6. Discussion
Firstly, due to computational resource constraints, the experiments are limited to models with a maximum size of 13B parameters. Larger models, such as those with 60B parameters or more, are not explored, which may exhibit different vulnerabilities to jailbreak attacks. Future studies could extend this research to larger-scale models to determine whether similar attack patterns and mitigation strategies apply.
@@
Secondly, while we select representative methods from established jailbreak taxonomy for analysis, some recent jailbreak methods are not included in this study. Nevertheless, our findings suggest that jailbreak methods within the same category share fundamentally similar jailbreak mechanisms, indicating the broad applicability of our research.
@@
Future direction. Building upon these insights, future work could design more robust and interpretable safety mechanisms for LLMs. Specifically, by leveraging the understanding of how jailbreak methods manipulate model components (e.g., key attention heads and MLP layers), we can develop more generalizable and resilient safeguards, which could target the specific areas of vulnerability identified, such as the suppression of refusal signals and the amplification of affirmation signals.
7. Conclusion
@@
In this paper, we propose JailbreakLens, a dual-perspective framework that explains the jailbreak mechanism from both representations and circuit levels. Through an in-depth analysis of seven jailbreak methods on the Llama and Vicuna models, we found jailbreak prompts amplify components that reinforce affirmative responses while suppressing those that produce refusal, which shifts model representations toward safe clusters to deceive the LLM, leading it to provide detailed responses instead of refusals. Moreover, we found a strong and general correlation between representation deception and circuit activation shift across different models and jailbreak methods. These insights deepen our understanding of how jailbreaks succeed and provide valuable guidance for developing stronger defense mechanisms.
Appendix A Supplementary of Experimental Setting
A.1. Architecture of the Probes
The probes we used are as follows:
A.2. Details of Jailbreak Attacks
Table 4. Attack success rate of harmful prompts and prompts with different jailbreak strategies.
@@ A.2. Details of Jailbreak Attacks
The attack success rate of the jailbreak strategies considered in this paper on the Advbench dataset is summarized in Tab. 4. Here are examples of each jailbreak method targeted Llama2-7b used for evaluation.
@@
Appendix B Supplementary Results of Circuit Analysis
The contributions to the refusal score of all attention heads in all models we analyzed are shown in Fig. 8.
For Vicuna1.5-7b, the location of affirmation and refusal signal attention heads is consistent with Llama2-7b, i.e., L26H4 and L21H14. For Llama2-13b and Vicuna1.5-13b, the location of affirmation and refusal signal attention heads is the same, i.e., L37H37 and L31H35. Moreover, the results of contributions to the refusal score of MLP layers are shown in Fig. 9.
The results of how MLP layers on Llama2-7b-chat-hf, Llama2-13b-chat-hf, Llama3-8b-Instrcution, Vicuna1.5-7b and Vicuna1.5-13b are activated with jailbreak prompts are summarized in Fig. 10, respectively.
Figure 8. Attention Heads attribution of refusal score in Llama2-7b, Llama2-13b, Llama3-8b, Vicuna1.5-7b, and Vicuna1.5-13b, respectively.
Figure 9. MLP layers attribution of refusal score on harmful and safe prompts in Llama2-7b, Llama2-13b, Llama3-8b, Vicuna1.5-7b, and Vicuna1.5-13b, respectively.
Figure 10. MLP layers attribution of refusal score on jailbreak prompts in Llama2-7b, Llama2-13b, Llama3-8b, Vicuna1.5-7b, and Vicuna1.5-13b, respectively.
Appendix C Supplementary Results of Evolution Analysis
@@
Figure 11. Dynamics of key components activation and representation on different jailbreak prompts in Llama2-7b.